Water Technology, Inc. Incurred the following costs during 20xt. The company sold all f ts products manufactured during the year Direct material Direct labor $5,000,000 2,400,000 Manufacturing overheadr Ueilities (primarily electricity) 120,000 220,000 150,000 400,000 230,000 Depreciation on plant and equipment Insurance Supervisory salaries Property taxes Selling costs Advertising Sales commissions Administrative costs Salaries of top management and staft 165,000 80,000 Office supplies Depreciation on building and equipment 372,000 45,000 80,000 to a slowdown in the economy. Prospects for 20x2 are slightly ces During 20x1, the company operated at about half of its capacity, due better. Jared Lowes, the marketing manager, forecasts a 20 percent growth in sales over the 20xt level. Required Categorize each of the costs listed below as to whether it is most likely variable or fixed. Forecast the 20x2 cost amount for each of the cost items listed Direct material Direct labor Manufacturing overhead Utlities (primarily electricity) Depreciation on plant and 19 of 19EENex Prex
Required: Categorize each of the costs listed below as to whether it is most likely variable or fixed. F the cost items listed. Variable or Fixed 20x2 Forecast Direct material Direct labor Manufacturing overhead Utilities (primarily electricity) Depreciation on plant and equipment Supervisory salaries Property taxes Selling costs: Advertising Sales commissions Administrative costs Salaries of top management and staff Office supplies Depreciation on building and equipment <Prey 19 of 19 Next F5 F7 F2 F3
Answer #1
Variable or fixed 20x2 Forecast
Direct material Variable 6000000 =5000000*120%
Direct labor Variable 2880000 =2400000*120%
Manufacturing overhead:
Utilities (primarily electricity) Variable 144000 =120000*120%
Depreciation on plant and equipment Fixed 220000
Insurance Fixed 150000
Supervisory salaries Fixed 400000
Property taxes Fixed 230000
Selling costs:
Advertising Fixed 165000
Sales commissions Variable 96000 =80000*120%
Administrative costs:
Salaries of top management and staff Fixed 372000
Office supplies Fixed 45000
Depreciation on building and equipment Fixed 80000
Note: Office Supplies can alternatively be categorized as Variable with amount of $54000